# Differentially Private Bilevel Optimization

### v2

@@ Appendix B Auxiliary Lemma
Plugging $\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})$ into the first inequality above, and $\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})$ into the second, shows that
00 0 $\leq\left\langle\mathbf{v}+\frac{1}{\eta}(\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathbf{x}),\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\rangle~{},$
00 0 $\leq\left\langle\mathbf{w}+\frac{1}{\eta}(\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathbf{x}),\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})\right\rangle=\left\langle-\mathbf{w}+\frac{1}{\eta}(\mathbf{x}-\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})),\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\rangle~{}.$
@@
00 0 $\leq\left\langle{\mathbf{v}-\mathbf{w}}+\frac{1}{\eta}(\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})),\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\rangle$
$=\left\langle\mathbf{v}-\mathbf{u},\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\rangle-\frac{1}{\eta}\left\|\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\|^{2}$
$\leq\left\|\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\|\left(\left\|\mathbf{v}-\mathbf{u}\right\|-\frac{1}{\eta}\left\|\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\|\right)~{}.$
@@
$\left\|\mathbf{v}-\mathbf{u}\right\|$ $\geq\frac{1}{\eta}\left\|\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})\right\|$

### v3

@@ Appendix C Auxiliary lemma
Plugging $\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})$ as $\mathbf{u}$ into the first inequality above, and $\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})$ into the second, shows that
$0$ $\leq\left\langle\mathbf{v}+\frac{1}{\eta}(\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathbf{x}),\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\rangle~,$
$0$ $\leq\left\langle\mathbf{w}+\frac{1}{\eta}(\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathbf{x}),\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})\right\rangle=\left\langle-\mathbf{w}+\frac{1}{\eta}(\mathbf{x}-\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})),\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\rangle~.$
@@
$0$ $\leq\left\langle{\mathbf{v}-\mathbf{w}}+\frac{1}{\eta}(\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})),\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\rangle$
$=\left\langle\mathbf{v}-\mathbf{w},\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\rangle-\frac{1}{\eta}\left\|\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\|^{2}$
$\leq\left\|\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\|\left(\left\|\mathbf{v}-\mathbf{w}\right\|-\frac{1}{\eta}\left\|\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\|\right)~.$
@@
$\left\|\mathbf{v}-\mathbf{w}\right\|$ $\geq\frac{1}{\eta}\left\|\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})\right\|$